OpenStudy (anonymous):

In a basketball tournament, team A scored 6 more points than 3 times as many points as team B scored. Team C scored 45 more points than team B scored. The combined scores for all three teams was 476 points. Which statement about the number of points is true? Team A scored 85 points. Team A scored 255 points. Team C scored 261 points. Team C scored 130 points.

OpenStudy (agl202):

A = 3B + 6 C = B + 45 A + B + C = 476 (3B + 6) + B + (B + 45) = 476 5B + 51 = 476 5B = 476 - 51 5B = 425 B = 85 A = 3B + 6 A = 3(85) + 6 A = 255 + 6 A = 261 C = B + 45 C = 85 + 45 C = 130 check... A + B + C = 476 261 + 85 + 130 = 476 476 = 476 (correct) A = 261, B = 85, C = 130 So correct answer is D. To give me a medal click the button to the right of my answer: Best Response------------------------------>